# Application of the FRADO model of BLR formation to the Seyfert galaxy NGC 5548 and the first step toward determining the Hubble constant

### v2

@@ 6.1 Covering factor
Although we refer to our disk-plus-BLR model as self-consistent, not all of its components are fully self-consistent yet. In principle, for a given black hole mass and Eddington ratio, the BLR covering factor should be determined by the model. Some tests of the predicted covering factor have been performed for relatively high Eddington ratios (Naddaf et al., 2023; Naddaf & Czerny, 2024). However, at lower Eddington ratios, the clouds are located too close to the disk, leading to inconsistencies.
@@ 6.2 FRADO vs LOC models as a method to reproduce the BLR transfer function
FRADO is a dynamical model which aims to reproduce the physical origin of the clouds, their location and their velocities. The model is not yet fully matured, but there are clear prospects for the model further development. For the dynamics, the addition of the line driving, mentioned above, should be included as its importance is increasing with the height of the clouds. Next, the emissivity of each clouds should be calculated individually, according to its location. Finally, there are two even more complex problems. One is a good realistic description of the shielding of clouds by themselves, important at low heights from the disk. The second one is determination of the density distribution of material based on thermal instability of the medius, as described already by Krolik et al. (1981). The wind is launched as a continuous medium, with clouds forming out of this wind in short thermal timescale.
LOC models, on the other hand, by their concept (power law distribution of clouds in radius and in density, with the limiting densities and limiting radii, and power law slopes as parameters) are better suitable for accurate matching of the line ratios in the AGN spectra. The additional parameters are still required if the distribution is flat, concentrated around the accretion disk, instead of a spherical distribution as in the original model of Baldwin et al. (1995). This model is mature, but its parameters must be controlled by the data. However, to derive the BLR response, additional assumptions are required. Lawther et al. (2018) combined spherically-symmetric approach with continuity equation to set the clod radial distribution but the authors indeed were able to determine the expected delay for all emission lines. Korista & Goad (2019), also using LOC model, pioneered the study of the disk continuum contamination by BLR continuum. However, it still remains to be seen if LOC models can be used in the same way as FRADO for fitting both the spectra and time delays, and for the Hubble constant determination. This was not done so far.
@@ 6.3 Other distance measurements to the source
The heliocentric redshift of NGC 5548 (; de Vaucouleurs et al., 1991) corresponds to a distance of Mpc when the motion of the source is measured relative to the Cosmic Microwave Background (CMB)222https://ned.ipac.caltech.edu/. The proper motion of the source is not negligible, as indicated by the difference between its heliocentric and CMB-based velocities reported by NED ( km s-1 vs. km s-1).
@@
Applying a velocity correction from the heliocentric frame to the CMB frame shifts the redshift from 0.01651 to 0.01721. This adjustment would, in turn, modify the our estimated value of the Hubble constant to km s-1 Mpc-1. However, given the relatively large uncertainties involved, achieving a highly precise correction for the proper motion is not crucial in this context.
@@ 6.4 The Hubble tension
In response to this discrepancy, several independent methods have been proposed to confirm or address the tension (e.g., see de Jaeger et al., 2022; Wong et al., 2024; Trefoloni et al., 2024; Liu et al., 2024; Bargiacchi et al., 2025, for some of the latest examples). Notably, some local measurements yield lower values of (e.g., Grillo et al., 2024). The latest findings from the clustering of galaxies, quasars, and Lyman- forest tracers, based on the first year of observations from the Dark Energy Spectroscopic Instrument (DESI Data Release 1; DESI Collaboration et al. 2024), support a lower Hubble constant of km s-1 Mpc-1 for a flat CDM model, with remarkably small uncertainty. Their results, incorporating constraints from Big Bang nucleosynthesis and acoustic angular scale measurements from the CMB, suggest km s-1 Mpc-1, while a joint analysis of CMB anisotropies and CMB lensing from Planck+ACT yields km s-1 Mpc-1.
@@ 6.5 Prospects for continuum reverberation from LSST
The standard six-color photometry provided by the Legacy Survey of Space and Time (LSST), soon to be available from the Vera Rubin Observatory, may not achieve the same level of precision in time delay measurements as combined SWIFT/ground-based monitoring. However, there remain promising opportunities for detecting continuum time delays, particularly for systems with larger black hole masses with high luminosities (Pozo Nuñez et al., 2023, 2024; Panda et al., 2024).
For the most massive quasars, however, the contribution of the BLR to the observed time delay may be less significant. This is suggested by results for , a quasar with a black hole mass of , as reported by Kokubo (2018). To better understand this effect, we can also investigate a potential scaling relation between BLR continuum contamination and emission line intensities, which may provide further insight into the role of the BLR in time delay measurements.
@@ 7 Conclusions
Our pilot study aimed to test the FRADO model for the BLR structure and estimate the Hubble constant based on the mean spectrum data and continuum time delay measurements for NGC 5548. The study has been highly successful, yielding several key outcomes:
@@
Relaxing the assumption of a known redshift-based luminosity, we model the spectrum and time-delay data across a range of Eddington ratios. For each value, we optimized the source distance, ultimately identifying a best-fit Eddington ratio and corresponding distance, which in turn provided an estimate of the Hubble constant.
Although the uncertainty in the Hubble constant remains relatively large ( km s-1 Mpc-1), our method demonstrates potential for refinement. Further improvements—both for this source and for similar determinations using other AGN with dense monitoring—could lead to significantly more precise measurements.
@@ Acknowledgements.
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. [951549]). VKJ acknowledges the OPUS-LAP/GA ˇCR-LA bilateral project (2021/43/I/ST9/01352/OPUS 22 and GF23-04053L). MHN also acknowledges the financial support by the University of Liege under Special Funds for Research, IPD-STEMA Program. SP is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. BC acknowledges the support from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). FPN gratefully acknowledges the generous and invaluable support of the Klaus Tschira Foundation.References
@@
- Chatzikos et al. (2023) Chatzikos, M., Bianchi, S., Camilloni, F., et al. 2023, Rev. Mexicana Astron. Astrofis., 59, 327
@@
- Kammoun et al. (2024) Kammoun, E., Papadakis, I. E., Dovčiak, M., & Panagiotou, C. 2024, A&A, 686, A69
@@
- Krolik (1999) Krolik, J. H. 1999, Active galactic nuclei : from the central black hole to the galactic environment
@@
- Pandey et al. (2024) Pandey, A., Martínez-Aldama, M. L., Czerny, B., Panda, S., & Zajaček, M. 2024, arXiv e-prints, arXiv:2401.18052

### v3

@@ 6.1 Covering factor
Although we refer to our disk-plus-BLR model as self-consistent, not all of its components are fully self-consistent yet. In principle, for a given and Eddington ratio, the BLR covering factor should be determined by the model. Some tests of the predicted covering factor have been performed for relatively high Eddington ratios (Naddaf et al., 2023; Naddaf & Czerny, 2024). However, at lower Eddington ratios, the clouds are located too close to the disk, leading to inconsistencies.
@@
6.2 Relativistic effects and X-ray albedo
Currently our model does not account for the effects of GR, which are critically important for phenomena within the inner 10 . In these inner regions, spacetime curvature significantly influences both the propagation of photons from the lamppost (a point-like X-ray source) to the accretion disk and their subsequent travel from the disk to the observer. Additionally, the energy dissipation profile in the cold disk depends on the black hole spin (see e.g., Novikov & Thorne, 1973).
However, the lamppost geometry itself is a considerable simplification, and in the case of NGC 5548, the cold accretion disk does not extend all the way to the ISCO. As a result, the reprocessing primarily occurs at larger radii, where GR effects are less dominant.
Another important consideration is the disk albedo, which is a complex function of the density in the disk atmosphere and the ionization parameter, both of which are likely to vary with radius. This issue is best addressed in systems that exhibit broad, relativistically smeared X-ray Fe K lines, where model assumptions can be more directly tested (e.g., Ballantyne et al., 2001; Dovčiak et al., 2004; Garc´ıa et al., 2013). Even models of reprocessing by the warm corona were studied recently by Petrucci et al. (2020).
For the case of continuum time delays, GR effects have been specifically studied for NGC 5548 by Kammoun et al. (2024b). However, these models do not include contributions from the BLR and fail to reproduce features such as the Balmer and Paschen edges in the time-delay spectrum.
In the future, it may become possible to integrate these various components into a single model. However, full GR-based computations are computationally expensive and require the inclusion of additional free functions (radial profile of the ionization parameter, disk geometrical height, hard X-ray source geometry) to model the reprocessing accurately. Overall, as demonstrated by Kammoun et al. (2021a), the disk response is indeed sensitive to both relativistic effects and albedo. Nonetheless, those models do not incorporate an inner warm corona and therefore cannot be used for direct comparison in our case.
6.3 Dust sublimation temperature
In this work, we adopt a universal dust sublimation temperature of 1500 K, which plays a crucial role in the FRADO model as it sets the onset of BLR cloud launching. This choice directly influences the analytical expression for the Hubble constant derived in equation 11. Given the temperature profile in the standard accretion disk that underlies this formula, our inferred values for the Hubble constant scale approximately as:
However, the dust sublimation temperature is not firmly established. While Barvainis (1987) originally recommended 1500 K as a representative value, more recent papers recommend 1500 K for standard mixture of silicate/graphite dust (e.g. Mor & Netzer, 2012), but for pure graphite, higher values such as 1800 K are proposed (e.g., Mor & Netzer, 2012). A broader, density-dependent range was presented by Baskin & Laor (2018), whose Figure 1 shows that at densities around , sublimation temperatures for silicate and graphite grains could reach 1600 K and 2000 K, respectively.
Nonetheless, graphite is unlikely to be present in the nuclear parts of AGN, as the 2175 Å feature is generally not seen in absorption (Laor & Draine, 1993; Gaskell et al., 2004), which rather suggests the presence of amorphous carbon grains (e.g. Czerny et al., 2004). However, there are no reliable estimates for the sublimation temperature of amorphous carbon in AGN environments. Stellar measurements from AGB stars give a temperature of K, very close to the sublimation temperature of silicate in these surroundings.
An alternative constraint was proposed by Vazquez et al. (2015), who argued for dust temperatures as high as 1800 K based on dust time-delay measurements rather than direct spectral analysis. However, their interpretation may be affected by an incomplete subtraction of the accretion disk contribution to the observed dust-delays (Mandal et al., 2024).
Given these uncertainties, we adopt 1500 K as a conservative and widely used value, but we acknowledge a plausible error margin of K. Based on the scaling relation above, this translates into a uncertainty in the inferred Hubble constant, representing an important systematic contribution to our results.
The presence of dust, particularly of dust with atypical properties, may additionally imply problems with establishing the intrinsic SED. For example, Gaskell et al. (2023) argue that, based on their extinction curve, the UV flux of NGC 5548 may be underestimated by a factor of seven. Tests of such issues can be performed in the future.
6.4 Other fixed parameters as sources of systematic errors
It would be computationally very time consuming to move the other fixed parameters to the parameters fitted to the data, and to study the degeneracy between all these parameters. However, we can easily discuss their expected importance looking at the analytical formula for the Hubble constant of Collier et al. (1999) (see equation 11). Those extra fixed parameters are: , inclination angle, gas density, and column density. The black hole mass does not enter into equation 11. The monochromatic flux is not sensitive on as long as the optical spectrum does not show strong departure from a power law (Shakura & Sunyaev, 1973). This happens only in the far UV. The presence of the warm corona complicates the reasoning. Also the shape of the structure function shows some dependence on (see e.g. Naddaf et al., 2025, for most recent study), and not just on local monochromatic flux which is actually fitted. The onset of the BLR does not depend on , as it is set by the dust sublimation temperature, but the vertical velocities increase with . Overall, it should not lead to significant effect. The viewing angle is important, as it is present in equation 11. For the viewing angle of 30 degrees, the Hubble constant would be lower by 6 %, and for 50 degrees it would be higher by 8 %. Therefore, this effect is important for any specific source. On the other hand, if the method is later applied to a large sample of objects at different redshifts, the viewing angle effects will average out, as there is no strong evidence for a redshift-dependent trend in the viewing angle or the dusty torus opening angle (e.g. Prince et al., 2022; Rałowski et al., 2024). However, assessing the influence of the local density and column density of the BLR clouds is considerably more challenging, as both parameters impact the ionization state and the effective optical depth of the region. These quantities can be effectively constrained by analyzing the shape of the delay enhancement near the Balmer and Paschen edges. Moreover, any reduction in effective optical depth can be compensated for by adjusting the covering factor, which is treated as a free parameter in the model. We certainly plan to refine and improve the model in future work.
@@ 6.5 FRADO vs LOC models as a method to reproduce the BLR transfer function
FRADO is a dynamical model which aims to reproduce the physical origin of the clouds, their location and their velocities. The model is not yet fully matured, but there are clear prospects for the model further development. For the dynamics, the addition of the line driving, mentioned above, should be included as its importance is increasing with the height of the clouds. Next, the emissivity of each clouds should be calculated individually, according to its location. Finally, there are two even more complex problems. One is a good realistic description of the shielding of clouds by themselves, important at low heights from the disk. The second one is determination of the density distribution of material based on thermal instability of the medium, as described already by Krolik et al. (1981). The wind is launched as a continuous medium, with clouds forming out of this wind in short thermal timescale.
Locally Optimally Emitting Cloud (LOC) models, on the other hand, by their concept (power law distribution of clouds in radius and in density, with the limiting densities and limiting radii, and power law slopes as parameters) are better suited for accurately reproducing the line ratios observed in AGN spectra. The additional parameters are still required if the distribution is flat, concentrated around the accretion disk, instead of a spherical distribution as in the original model of Baldwin et al. (1995). This model is mature, but its parameters must be controlled by the data. However, to derive the BLR response, additional assumptions are required. Lawther et al. (2018) combined a spherically symmetric approach with the continuity equation to constrain the radial distribution of clouds, and were thereby able to predict the expected time delays for all emission lines. Additionally, using LOC model, Korista & Goad (2019) pioneered the study of the disk continuum contamination by BLR continuum. However, it remains to be seen whether LOC models can be applied in the same way as FRADO to simultaneously fit both the spectra and time delays, and thereby enable determination of the Hubble constant. This has not been achieved so far.
@@ 6.6 Other distance measurements to the source
The heliocentric redshift of NGC 5548 (; de Vaucouleurs et al., 1991) corresponds to a distance of Mpc. It is different when the motion of the source is measured relative to the Cosmic Microwave Background (CMB)222https://ned.ipac.caltech.edu/. The proper motion of the source is thus not negligible, as indicated by the difference between its heliocentric and CMB-based velocities reported by NED ( km s-1 vs. km s-1).
@@
Applying a velocity correction from the heliocentric frame to the CMB frame shifts the redshift from 0.017175 to 0.017876. This adjustment would, in turn, modify our estimated value of the Hubble constant to km s-1 Mpc-1. However, given the relatively large uncertainties involved, achieving a highly precise correction for the proper motion is not crucial in this context.
@@ 6.7 The Hubble tension
In response to this discrepancy, several independent methods have been proposed to confirm or address the tension (e.g., see de Jaeger et al., 2022; Wong et al., 2024; Trefoloni et al., 2024; Liu et al., 2024; Bargiacchi et al., 2025, for some of the latest examples). Notably, some local measurements yield lower values of (e.g., Grillo et al., 2024). The latest findings from the clustering of galaxies, quasars, and Lyman- forest tracers, based on the first year of observations from the Dark Energy Spectroscopic Instrument (DESI Data Release 1; DESI Collaboration et al. 2024), support a lower Hubble constant of km s-1 Mpc-1 for a flat CDM model, with remarkably small uncertainty. Their results, incorporating constraints from Big Bang nucleosynthesis and acoustic angular scale measurements from the CMB, suggest km s-1 Mpc-1, while a joint analysis of CMB anisotropies and CMB lensing from Planck+ACT yields km s-1 Mpc-1. Most recent measurements from the Dark Energy Survey (DES) experiment yields an value of km s-1 Mpc-1 (DES Collaboration et al., 2025).
@@ 6.8 Prospects for continuum reverberation from LSST
The standard six-color photometry provided by the Legacy Survey of Space and Time (LSST), soon to be available from the Vera Rubin Observatory, may not achieve the same level of precision in time delay measurements as combined SWIFT/ground-based monitoring. However, there remain promising opportunities for detecting continuum time delays, particularly for AGNs with larger black hole masses with high luminosities (Pozo Nuñez et al., 2023, 2024; Panda et al., 2024).
For the most massive quasars, however, the contribution of the BLR to the observed time delay may be less significant. This is suggested by results for , a quasar with of , as reported by Kokubo (2018). To better understand this effect, we can also investigate a potential scaling relation between BLR continuum contamination and emission line intensities, which may provide further insight into the role of the BLR in time delay measurements.
@@ 7 Conclusions
Our pilot study aims to test the FRADO model for the BLR structure and estimate the Hubble constant based on the mean spectrum data and continuum time delay measurements for NGC 5548. The study has been highly successful, yielding several key outcomes:
@@
Relaxing the assumption of a known redshift-based luminosity, we model the spectrum and time-delay data across a range of Eddington ratios. For each value, we optimize the source distance, ultimately identifying a best-fit Eddington ratio and corresponding distance, which in turn provide an estimate of the Hubble constant.
Although the uncertainty in the Hubble constant remains relatively large ( km s-1 Mpc-1), our method demonstrates potential for refinement. Further improvements, both for this source and for similar determinations using other AGNs with dense monitoring could lead to significantly more precise measurements.
@@ Acknowledgements.
We thank the anonymous referee for valuable comments and suggestions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. [951549]). VKJ acknowledges the OPUS-LAP/GA ˇCR-LA bilateral project (2021/43/I/ST9/01352/OPUS 22 and GF23-04053L). MHN also acknowledges the financial support by the University of Liege under Special Funds for Research, IPD-STEMA Program. SP is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. BC acknowledges the support from COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). FPN gratefully acknowledges the generous and invaluable support of the Klaus Tschira Foundation. We thank Dr. Missagh Mehdipour for sharing the broadband SED data points with us.References
@@
- Barvainis (1987) Barvainis, R. 1987, ApJ, 320, 537
- Baskin & Laor (2018) Baskin, A. & Laor, A. 2018, MNRAS, 474, 1970
@@
- Cao et al. (2025) Cao, S., Mandal, A. K., Zajaček, M., Czerny, B., & Ratra, B. 2025, Phys. Rev. D, 111, 083545
- Cao et al. (2024) Cao, S., Zajaček, M., Czerny, B., Panda, S., & Ratra, B. 2024, MNRAS, 528, 6444
- Cao et al. (2022) Cao, S., Zajaček, M., Panda, S., et al. 2022, MNRAS, 516, 1721
- Cappi et al. (2016) Cappi, M., De Marco, B., Ponti, G., et al. 2016, A&A, 592, A27
- Chatzikos et al. (2023a) Chatzikos, M., Bianchi, S., Camilloni, F., et al. 2023a, Rev. Mexicana Astron. Astrofis., 59, 327
- Chatzikos et al. (2023b) Chatzikos, M., Bianchi, S., Camilloni, F., et al. 2023b, Rev. Mexicana Astron. Astrofis., 59, 327
@@
- Crenshaw et al. (2009) Crenshaw, D. M., Kraemer, S. B., Schmitt, H. R., et al. 2009, ApJ, 698, 281
- Czerny et al. (2018) Czerny, B., Beaton, R., Bejger, M., et al. 2018, Space Sci. Rev., 214, 32
@@
- Dalla Bontà et al. (2020) Dalla Bontà, E., Peterson, B. M., Bentz, M. C., et al. 2020, ApJ, 903, 112
@@
- DES Collaboration et al. (2025) DES Collaboration, Abbott, T. M. C., Acevedo, M., et al. 2025, arXiv e-prints, arXiv:2503.06712
@@
- Ferland et al. (1979) Ferland, G. J., Netzer, H., & Shields, G. A. 1979, ApJ, 232, 382
- Floris et al. (2025) Floris, A., Pandey, A., Czerny, B., et al. 2025, A&A, 697, A23
@@
- Garc´ıa et al. (2013) García, J., Dauser, T., Reynolds, C. S., et al. 2013, ApJ, 768, 146
- Gaskell et al. (2023) Gaskell, C. M., Anderson, F. C., Birmingham, S. Á., & Ghosh, S. 2023, MNRAS, 519, 4082
- Gaskell et al. (2004) Gaskell, C. M., Goosmann, R. W., Antonucci, R. R. J., & Whysong, D. H. 2004, ApJ, 616, 147
@@
- Kammoun et al. (2024a) Kammoun, E., Papadakis, I. E., Dovčiak, M., & Panagiotou, C. 2024a, A&A, 686, A69
- Kammoun et al. (2024b) Kammoun, E., Papadakis, I. E., Dovčiak, M., & Panagiotou, C. 2024b, A&A, 686, A69
@@
- Krolik (1999a) Krolik, J. H. 1999a, Active galactic nuclei : from the central black hole to the galactic environment
- Krolik (1999b) Krolik, J. H. 1999b, Active galactic nuclei : from the central black hole to the galactic environment
@@
- Li et al. (2025) Li, Y.-R., Shangguan, J., Wang, J.-M., et al. 2025, arXiv e-prints, arXiv:2502.18856
- Li et al. (2013) Li, Y.-R., Wang, J.-M., Ho, L. C., Du, P., & Bai, J.-M. 2013, ApJ, 779, 110
- Li et al. (2022) Li, Y.-R., Wang, J.-M., Songsheng, Y.-Y., et al. 2022, ApJ, 927, 58
@@
- Lu et al. (2016) Lu, K.-X., Du, P., Hu, C., et al. 2016, ApJ, 827, 118
- Lusso et al. (2025) Lusso, E., Risaliti, G., & Nardini, E. 2025, arXiv e-prints, arXiv:2504.02040
- Mandal et al. (2025) Mandal, A. K., Woo, J.-H., & Wang, S. 2025, ApJ, 985, 30
- Mandal et al. (2024) Mandal, A. K., Woo, J.-H., Wang, S., et al. 2024, ApJ, 968, 59
- Mart´ınez-Aldama et al. (2019) Martínez-Aldama, M. L., Czerny, B., Kawka, D., et al. 2019, ApJ, 883, 170
@@
- Naddaf et al. (2025) Naddaf, M. H., Martinez-Aldama, M. L., Hutsemekers, D., Savic, D., & Czerny, B. 2025, arXiv e-prints, arXiv:2506.01159
@@
- Novikov & Thorne (1973) Novikov, I. D. & Thorne, K. S. 1973, in Black Holes (Les Astres Occlus), ed. C. Dewitt & B. S. Dewitt, 343–450
- Pancoast et al. (2014) Pancoast, A., Brewer, B. J., & Treu, T. 2014, MNRAS, 445, 3055
@@
- Pandey et al. (2025) Pandey, A., Martínez-Aldama, M. L., Czerny, B., et al. 2025, ApJS, 277, 36
@@
- Prince et al. (2025) Prince, R., Hernández Santisteban, J. V., Horne, K., et al. 2025, arXiv e-prints, arXiv:2506.06731
- Prince et al. (2022) Prince, R., Hryniewicz, K., Panda, S., Czerny, B., & Pollo, A. 2022, ApJ, 925, 215
@@
- Rosborough et al. (2025) Rosborough, S., Robinson, A., Almeyda, T., Humphrey, D., & Noll, M. 2025, arXiv e-prints, arXiv:2507.14337
- Rosborough et al. (2024) Rosborough, S. A., Robinson, A., Almeyda, T., & Noll, M. 2024, ApJ, 965, 35
@@
- Shimura & Takahara (1995) Shimura, T. & Takahara, F. 1995, ApJ, 445, 780
- Śniegowska et al. (2021) Śniegowska, M., Marziani, P., Czerny, B., et al. 2021, ApJ, 910, 115
@@
- Suyu et al. (2024) Suyu, S. H., Goobar, A., Collett, T., More, A., & Vernardos, G. 2024, Space Sci. Rev., 220, 13
@@
- Wildy et al. (2021) Wildy, C., Landt, H., Ward, M. J., Czerny, B., & Kynoch, D. 2021, MNRAS, 500, 2063
@@
- Zajaček et al. (2024) Zajaček, M., Panda, S., Pandey, A., et al. 2024, A&A, 683, A140
Appendix A based on observationally determined response function
We employ an alternative method to validate the results obtained from the FRADO model. In rare cases, NGC 5548 being one notable example, the transfer function for H was directly determined from the data using densely sampled spectroscopic RM. For this purpose, we use the results presented by Horne et al. (2021). These observations were conducted in 2014, roughly during the same period with the photometric time-delay measurements. The study provides response functions for various emission lines, but we focus on those corresponding to the H line, as they are most representative of the regions in the BLR where the Balmer and Paschen continua originate and play a dominant role in shaping the observed time delays. Specifically, we consider two versions of the H response functions presented by Horne et al. (2021): one that encompasses the full H line and another centered narrowly around 4863 Å. A comparison between our model-derived BLR response function and the two response functions from Horne et al. (2021) is shown in the upper panel of Figure 12.
Both observed response functions lack a double-peak structure, which, however, is present in the case of the CIV line (see Figure 4 of Horne et al. 2021). More importantly, these response functions exhibit a significantly longer tail, with this extended tail being notably stronger in the full-H response profile than in the narrower 4863 Å case. This long tail is also clearly visible in the 4830 Å bin shown in Figure 9 of Horne et al. (2021), suggesting that the extended delay may arise from contamination by Fe II emission. On average, the optical and UV Fe II delays tend to be longer than those of typical low-ionization lines such as Mg II and H (Zajaček et al. 2024).
Having an observationally derived transfer function significantly reduces our dependence on the FRADO model summarized in Section 3.3. However, we still require CLOUDY computations to determine the wavelength-dependent efficiency of re-emission by the BLR, as well as a disk/corona model.
Therefore, we perform fitting of both the mean spectrum and the wavelength-dependent time delay as before, for a set of values of the luminosity distance. For each luminosity distance , the model parameters are refitted accordingly.
In the case of the entire H transfer function, we are unable to tightly constrain the luminosity distance (blue data points in the bottom panel of 12); the monotonic trend continues up to 98 Mpc, with marginal rise at 102 Mpc. We did not perform computations beyond this distance. However, for the transfer function centered around 4863 Å , a broad minimum is observed from 82 Mpc till 94 Mpc (red data points), which is shifted compared to the 74 Mpc value obtained using the theoretical FRADO model for the transfer function (black data points). Interestingly, the overall fit is the best for the FRADO model. The resulting constraint on the Hubble constant from the 4863 Å bin is km s-1 Mpc-1, generally lower than the FRADO-derived range but only marginally outside the 1 sigma error value due to relatively large uncertainty, further supporting the applicability of the FRADO model in such analyses.